Harry Hargrove invented a sequence but he did not tell what method he used. What is the 10th term of his sequence 2,6,24,120?
i need the answer asap ill appreciate it very much.

Answer by greenestamps(13200)



It's unfortunate that you need the answer "ASAP"... because it is IMPOSSIBLE to know the answer to the question.

The same is true of ANY problem like this where a sequence of numbers is given and you are supposed to find the next numbers, or missing numbers.

Given ANY problem like this, you can put any numbers you want in the missing spots and it will be a valid sequence.

So, technically, the question is not valid; you can't KNOW "THE" answer using formal mathematics.

On the other hand, there is an obvious pattern for the numbers in this example which make it VERY LIKELY what the next numbers are:
2*3 = 6
6*4 = 24
24*5 = 120
...

In fact, the given numbers are

2!, 3!, 4!, 5!, ...

So the LIKELY answer to the problem is that the 10th term of the sequence is 11!.

But there is NO WAY to KNOW if that is right....